OpenStudy (anonymous):

Please Help! 1/8x^2+6 Find concave up and concave down intervals also inflection point

OpenStudy (anonymous):

I found f'=-16x/(8x^2+6)^2 f'' not sure if this is right but i got (-1024x^4+2560x^2+2496)/(8x^2+6)^4

OpenStudy (campbell_st):

well thats ok.... now solve f'(x) = 0 this will only occur when x = 0 that's the only stationary point test it in the 2nd derivative and you get and I think there is an error in the 2nd derivative I found \[f'(x) = -16x(8x^2 +6)^{-2}\] and before simplifying... I used the product rule rather then quotient \[f''(x) = \frac{512x^2 -128x - 96}{(8x^2 + 6)^3}\] so testing the stationary point x = 0 f''(0) = -96/6^3 less than zero so x = 0 is a max... the denominator will always be positive so no vertical asymptotes and a horizontal asymptote at x = 0 substitute it into the original equation and you'll find that the max value is 1/6 hope it helps

OpenStudy (campbell_st):

so since you only have a max... it concave down... I suppose.. it looks like